OpenStudy (anonymous):

How to factor \[3x^{\frac{3}{2}} - 9x^{\frac{1}{2}} + 6x^{\frac{-1}{2}}\]

OpenStudy (anonymous):

rfresh it

Parth (parthkohli):

Convert them into improper fractions first.

OpenStudy (anonymous):

improper fractions?

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow a{b \over c} = {ac + b \over c} }\)

OpenStudy (callisto):

\[3x^{\frac{3}{2}}-9x^{\frac{1}{2}}+6x^{-\frac{1}{2}}\]\[=3x^{-\frac{1}{2}}(x^{\frac{3}{2}+\frac{1}{2}}-3x^{\frac{1}{2}+\frac{1}{2}}+2)\]\[=3x^{-\frac{1}{2}}(x^2-3x+2)\]\[=3x^{-\frac{1}{2}}(x-2)(x-1)\]

Parth (parthkohli):

Oh wait, these were exponents?

OpenStudy (anonymous):

i dont get what @Callisto did either :/

OpenStudy (callisto):

Take out x^{-1/2} as a common factor, then factorise it

OpenStudy (anonymous):

uhmm can we like use quadratics or something? i dont know..you're better at me with this :(

OpenStudy (callisto):

Quadratic? yes when you factorise (x^2 -3x +2) But do it by factoring the expression is easier... and answer is trivial :|

OpenStudy (anonymous):

i can see :/ pls guide me

OpenStudy (callisto):

if you can see, why need I guide? (no offense...)

OpenStudy (anonymous):

lol :p i can see what you mean but i dont know how to do this

OpenStudy (callisto):

.... Sorry.. x^2 -3x +2 = x^2 -2x -x +2 = x(x-2) - (x-2) = (x-1)(x-2)

OpenStudy (anonymous):

now you're just confusing me

OpenStudy (anonymous):

how did you get that x^2- 3x + 2 again?

OpenStudy (callisto):

...... Please refer to the second step.

OpenStudy (anonymous):

but how do you know what exponent to factor out? i mean is there a way to know?

OpenStudy (callisto):

Finding the HCF of the terms...

OpenStudy (anonymous):

i always get confused with factoring fractional exponents sorry @FoolForMath

OpenStudy (anonymous):

ohhhh it's x^3/2 - (-1/2) that right?

OpenStudy (callisto):

The problem is the HCF in this case is 3x^{-1/2}

OpenStudy (callisto):

x^[3/2 - (-1/2) ] is for the first term

OpenStudy (anonymous):

oohhh ok...carry on..so we do qaudratics?

OpenStudy (callisto):

Why quadratic?

OpenStudy (anonymous):

i dont know..you said we do after factoring

OpenStudy (callisto):

...... You were asking factorization.. So I did factorization. Using quadratic (formula) to do factorization is kind of cheating, though you can use it.

OpenStudy (callisto):

Welcome

OpenStudy (anonymous):

o_o iss it me or did your smartscore just shoot up? xD

OpenStudy (anonymous):

haha